# Black Hole Orbit Classification: The Synergistic Effects of Cloud Strings, Gauss-Bonnet Terms, and Non-Commutative Parameters in Identifying WGC candidate Models: WGC as WCCC protector

### v1

@@ Abstract
The integration of non-commutative geometry and Gauss-Bonnet corrections in an action and the study of their black hole responses can provide highly intriguing insights. Our primary motivation for this study is to understand the interplay of these two parameters on the geodesics of spacetime, including photon spheres and time-like orbits. In this study, we found that this integration, in its initial form, can limit the value of the Gauss-Bonnet parameter ($\alpha$), creating a critical threshold beyond which changes in the non-commutative parameter ($\Xi$) become ineffective, and the structure can only manifest as a naked singularity. Furthermore, we found that using a more complex model, which includes additional factors such as a cloud of strings and linear charge, as a sample for studying spacetime geodesics, yield different and varied results. In this scenario, negative $\alpha$ values can also play a role, notably preserving the black hole form even with a super-extremal charge ($q>m$). For $\alpha>0.1$, the black hole mass parameter becomes significantly influential, with a critical mass below which the impact of other parameter changes is nullified. Interestingly, considering a more massive black hole, this high-mass state also maintains its black hole form within the super-extremal charge range. The existence of these two models led us to our main goal. By examining the temperature for these two cases, we find that both situations are suitable for studying the WGC. Finally, based on the behavior of these two models, we will explain how the WGC acts as a logical solution and a protector for the WCCC.
Non-commutativity, Gauss-Bonnet, black holes, photon spheres, Weak Gravity Conjecture
@@ I Introduction
In the previous article 1 , we discussed that due to the theoretical and practical limitations of current human knowledge, it is currently impossible to achieve a unified and comprehensive model for the structure of black holes. Therefore, various theoretical models are constructed and examined based on the combination of existing and known fields, with the aim of selecting the model that best aligns with nature and empirical data in the future 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 ; 10 . One of these models utilizes the properties of Non-Commutativity(NC), means the space-time coordinates do not commute, to study the geometry of spacetime. This model, while attempting to resolve the central singularity, also impacts the physical properties of black holes such as the event horizon and mass distribution. Our first goal in this series of articles is to examine these impact on the structure of geodesics and circular orbits of black holes influenced by NC. In previous work 1 , we explored models resulting from the combination of gravity and electromagnetism in both linear and nonlinear forms. We observed that in the linear form, despite the mass distribution changing from a point to a Gaussian distribution, the structural behavior of geodesics, including photon spheres and Time-like Circular Orbits(TCOs), remained very close to the Reissner-Nordström form. However, this combination imposed a constraint on the NC parameter, $\Xi$, such that, for example, given the selected parameter values, the structure retained its black hole form only within the range $0<\Xi\leq 0.2184$. In the nonlinear form, we used a model incorporated the Born-Infeld field and added the cosmological constant to the model. We found that the inclusion of the nonlinear field significantly improved the model’s accuracy, such that in this case, the parameter $\Xi$, based on the chosen values, could only lead to a standard model within the range $0<\Xi\leq 0.0611$. Since smaller values of $\Xi$ result in better alignment with reality, this reduction in range is quite significant. In this model, we observed that unlike the AdS mode, in the dS model, the curvature constant played a much more critical role, creating a critical curvature constant beyond which the structure would no longer exhibit black hole behavior. Finally, we examined the possibility of the existence of super-extremal black holes for the two models and found that only the Born-Infeld model has the capacity to form super-extremal black holes. The importance of super-extremal black holes lies in the fact that in this state, the likelihood of the existence of super-extremal particles 1 is higher, which, in addition to reinforcing the WGC, also strengthens the Weak Cosmic Censorship Conjecture (WCCC). Since this somehow shows that, even black holes with superextreme charges can hide their central singularity behind the event horizon.
@@
In Section 2, since in the previous work 1 we have sufficiently covered the principles and foundations, we will refer to that and only highlight the main equations. In Sections 3 and 4, we will examine and analyze the various models introduced using the discussed methods. Then, in Section 5, in search of evidence for the WGC, we will address the temperature structure in the super-extremal state and the conditions that could strengthen the WGC. Finally, in Section 6, we will present the results of our analysis.
@@ II.1 Topological Photon Sphere
To study the photon sphere, we use the topological method instead of the traditional method, a different method that, due to the general characteristics of its mathematical structure 21.1 ; 21.2 , has been extended not only to the study of light rings but also to thermodynamics21.3 ; 21.4 ; 21.5 ; 21.6 ; 21.7 ; 21.8 ; 21.9 .
Now we consider a general vector field as $\phi$ which can be decomposed into two components, $\phi^{r}$ and $\phi^{\theta}$,18
$\phi=(\phi^{r},\phi^{\theta}),$ (1)
also here, we can rewrite the vector as $\phi=||\phi||e^{i\Theta}$, where $||\phi||=\sqrt{\phi^{a}\phi^{a}}$, or $\phi=\phi^{r}+i\phi^{\theta}$. Based on this, the normalized vector is defined as,
$n^{a}=\frac{\phi^{a}}{||\phi||},$ (2)
where $a=1,2$ and $(\phi^{1}=\phi^{r})$ , $(\phi^{2}=\phi^{\theta})$. In the next stage, with respect to the necessity of spherical symmetry as a prerequisite for studying this method, and given the most general form of the metric in 4-dimensional form we have:
@@
$\begin{split}H(r,\theta)=\sqrt{\frac{-g_{tt}}{g_{\varphi\varphi}}}=\frac{1}{\sin\theta}\bigg{(}\frac{f(r)}{h(r)}\bigg{)}^{1/2}.\end{split}$ (4)
Based on this potential, we now define the vector field vector field $\phi=(\phi^{r},\phi^{\theta})$ as follows:
$\begin{split}&\phi^{r}=\frac{\partial_{r}H}{\sqrt{g_{rr}}}=\sqrt{g(r)}\partial_{r}H,\\ &\phi^{\theta}=\frac{\partial_{\theta}H}{\sqrt{g_{\theta\theta}}}=\frac{\partial_{\theta}H}{\sqrt{h(r)}}.\end{split}$ (5)
Now,using the above relations and considering the definition of topological charges, each photon sphere can be assigned a charge. To avoid repetition, a detailed description of how to calculate it is available in references 1 ; 18 ; 19 ; 20 ; 21 .
@@ II.2 Time-like Circular Orbits (TCOs)
Beyond the event horizon, it appears that our field equations remain logically applicable, allowing us to precisely identify and classify these geometric-kinematic boundaries. Accordingly, we have:
1. Null Geodesics (Light-like): The paths of massless particles and photons.
2. Timelike Geodesics: The trajectories of massive objects.
These geodesics, depending on the characteristics of the gravitational potential model, can either lack observable turning points or possess a turning point, forming loops. In null geodesics, these loops are referred to as photon spheres, while in timelike geodesics, they are known as TCOs (Timelike Circular Orbits). Both types of orbits can be stable or unstable.
Our assumptions in this study include a static and axisymmetric spacetime with $\mathbb{Z}_{2}$ symmetry in a 1+3 dimensional framework. Given these symmetries, discussing the equatorial plane does not compromise the generality of the analysis. Based on the metric equation ,Eq. (3), We consider the following quantities 22 :
@@
which E,L are the energy and the angular momentum. Now the Lagrangian can be recast as:
$2\mathfrak{L}=-\frac{A}{B}=\zeta,$ (8)
where $\zeta=-1,0$ for time-like, null geodesics, respectively. Using the above Lagrangian, the effective potential can be rewritten as follows:
@@
The concept of angular velocity (as measured by an observer at infinity) and $\beta$ in terms of metric parameters.
@@ III Non-Commutative black hole in 4D Einstein-Gauss-Bonnet theory
where $\Xi$ is NC parameter, m is the total mass of the source,$\alpha$ is a coupling constant with dimension of $(length)^{2}$ which is positive in heterotic string theory and the $\gamma$ is the lower incomplete Gamma function which can be written in terms of $\Gamma$ (the upper incomplete Gamma function) in the following form:
@@ III.1 Topological Photon Sphere
$\begin{split}H=\frac{\sqrt{4+\frac{\left(1-\sqrt{h}\right)r^{2}}{\alpha}}}{2\sin\!\left(\theta\right)r}.\end{split}$ (16)
@@
$\begin{split}\phi^{r}=-\frac{\left(\varphi_{1}+\varphi_{2}\right)\csc\!\left(\theta\right)}{\pi^{\frac{1}{4}}\sqrt{\frac{16\sqrt{\pi}\,\mathrm{erf}\left(\frac{\sqrt{\frac{r^{2}}{\Xi}}}{2}\right)\alpha m+\sqrt{\pi}\,r^{3}-16\sqrt{\frac{r^{2}}{\Xi}}\,{\mathrm{e}}^{-\frac{r^{2}}{4\Xi}}\alpha m}{r^{3}}}\,r^{3}\Xi}.\end{split}$ (17)
$\begin{split}\phi^{\theta}=-\frac{\sqrt{4+\frac{\left(1-\sqrt{h}\right)r^{2}}{\alpha}}\,\cos\!\left(\theta\right)}{2\sin\!\left(\theta\right)^{2}r^{2}}.\end{split}$ (18)
@@
Figure 2: Fig (2a): The normal vector field $n$ in the $(r-\theta)$ plane. The photon sphere is located at $(r,\theta)=(2.736931688316,1.57)$ with respect to $(\Xi=0.001,\alpha=0.39,m=1)$, (2b): the topological potential H(r) for NCEGB BH
This photon sphere has a Total Topological Charge (TTC) of -1 Fig. (2) and, as shown in Fig. (2), represents an unstable maximum. Consequently, this case presents a black hole containing an unstable photon sphere.
@@
Figure 3: Fig (3a): The normal vector field $n$ in the $(r-\theta)$ plane. The photon spheres are located at $(r,\theta)=(2.285600868252,1.57)$ and $(r,\theta)=(2.305792105113,1.57)$ with respect to $(\Xi=0.2488,\alpha=0.39,m=1)$, (3b): the topological potential H(r) for NCEGB BH
With $\Xi=0.2488$, the model in this horizonless region exhibits two photon spheres with topological charges of -1 and 1, indicating a spacetime with a TTC of zero Fig. (3). From an energy perspective Fig. (3), this situation corresponds to the presence of both a minimum and a maximum. Under these conditions, the structure manifests as a naked singularity.
@@
Table 1: *Unauthorized region: is the region where the roots of $\phi$ equations become negative or imaginary. $R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.
@@
Table 2: *Unauthorized region: is the region where the roots of $\phi$ equations become negative or imaginary in this region. TTC: *Total Topological Charge
@@ III.2 TCOs
$A=E^{2}r^{2}+\frac{L^{2}\left(r^{2}\sqrt{\frac{16\sqrt{\pi}\,\mathrm{erf}\left(\frac{\sqrt{\frac{r^{2}}{\Xi}}}{2}\right)\alpha m+\sqrt{\pi}\,r^{3}-16{\mathrm{e}}^{-\frac{r^{2}}{4\Xi}}\sqrt{\frac{r^{2}}{\Xi}}\,\alpha m}{r^{3}}}-\left(r^{2}+4\alpha\right)\pi^{\frac{1}{4}}\right)}{4\pi^{\frac{1}{4}}\alpha}.$ (19)
$B=-\frac{r^{2}\left(r^{2}\sqrt{\frac{16\sqrt{\pi}\,\mathrm{erf}\left(\frac{\sqrt{\frac{r^{2}}{\Xi}}}{2}\right)\alpha m+\sqrt{\pi}\,r^{3}-16{\mathrm{e}}^{-\frac{r^{2}}{4\Xi}}\sqrt{\frac{r^{2}}{\Xi}}\,\alpha m}{r^{3}}}-\left(r^{2}+4\alpha\right)\pi^{\frac{1}{4}}\right)}{4\pi^{\frac{1}{4}}\alpha}.$ (20)
@@
Figure 4: Fig (4a): $\beta$ diagram for structure in black hole form, (4b): MSCO localization and space classification for the black hole mode
@@
Figure 5: Fig (5a): $\beta$ diagram for structure in naked singularity form, (5b): Enlarging on the forbidden area in the diagram $\beta$ (5c): MSCO localization and space classification for the naked singularity mode

### v2

# WGC as WCCC protector: The Synergistic Effects of various Parameters in Identifying WGC candidate Models

@@ Abstract
The integration of non-commutative geometry and Gauss-Bonnet corrections in an action and the study of their black hole responses can provide highly intriguing insights. Our primary motivation for this study is to understand the interplay of these two parameters on the geodesics of spacetime, including photon spheres and time-like orbits. In this study, we found that this integration, in its initial form, can limit the value of the Gauss-Bonnet parameter ($\alpha$), creating a critical threshold beyond which changes in the non-commutative parameter ($\Xi$) become ineffective, and the structure can only manifest as a naked singularity. Furthermore, we found that using a more complex model, which includes additional factors such as a cloud of strings and linear charge, as a sample for studying spacetime geodesics, yield different and varied results. In this scenario, negative $\alpha$ values can also play a role, notably preserving the black hole form even with a super-extremal charge ($q>m$). For $\alpha>0.1$, the black hole mass parameter becomes significantly influential, with a critical mass below which the impact of other parameter changes is nullified. Interestingly, considering a more massive black hole, this high-mass state also maintains its black hole form within the super-extremal charge range. The existence of these two models led us to our main goal. By examining the temperature for these two cases, we find that both situations are suitable for studying the Weak Gravity Conjecture (WGC). Finally, based on the behavior of these two models, we will explain how the WGC acts as a logical solution and a protector for the WCCC.
Non-commutativity, Gauss-Bonnet, Cloud of Strings, black holes, photon spheres, Weak Gravity Conjecture
@@ I Introduction
In the previous article 1 , we discussed that due to the theoretical and practical limitations of current human knowledge, it is currently impossible to achieve a unified and comprehensive model for the structure of black holes. Therefore, various theoretical models are constructed and examined based on the combination of existing and known fields, with the aim of selecting the model that best aligns with nature and empirical data in the future 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 ; 10 . One of these models utilizes the properties of Non-Commutativity(NC), means the space-time coordinates do not commute, to study the geometry of spacetime. This model, while attempting to resolve the central singularity, also impacts the physical properties of black holes such as the event horizon and mass distribution. Our first goal in this series of articles is to examine these impact on the structure of geodesics and circular orbits of black holes influenced by NC. In previous work 1 , we explored models resulting from the combination of gravity and electromagnetism in both linear and nonlinear forms. We observed that in the linear form of charge, despite the mass distribution changing from a point to a Gaussian distribution, the structural behavior of geodesics, including photon spheres and Time-like Circular Orbits(TCOs), remained very close to the Reissner-Nordström form. However, this combination imposed a constraint on the NC parameter, $\Xi$, such that, for example, given the selected parameter values, the structure retained its black hole form only within the range $0<\Xi\leq 0.2184$. In the nonlinear form, we used a model incorporated the Born-Infeld field and added the cosmological constant to the model. Since smaller values of $\Xi$ result in better alignment with reality, We found that the inclusion of the nonlinear field significantly improved the model’s accuracy. Also, we observed that, in the dS model, the curvature constant played a much more critical role, creating a critical curvature constant beyond which the structure would no longer exhibit black hole behavior.
The primary objective of this series of articles is to develop models that can maintain the necessary initial conditions for black hole formation in both extremal and super-extremal charge regimes. Specifically, these models should simultaneously exhibit an event horizon and an unstable photon sphere. The significance of super-extremal black holes lies in the increased likelihood of the existence of ultra-extremal particles. This not only strengthens the Weak Gravity Conjecture (WGC) but also, as we will demonstrate, reinforces the Weak Cosmic Censorship Conjecture (WCCC). In our previous study, we found that only the Born-Infeld model exhibits such capacity [1].
@@
In Section 2, since in the previous work 1 we have sufficiently covered the principles and foundations, we will refer to that and only highlight the main equations. In the next two sections, we will study different models using the mathematical method introduced. In Section 5, we examine extreme and superextreme conditions and show that adding the temperature condition to the models under consideration can be a complement to make these definitions more precise, and accordingly, we can better understand the conditions that will lead to the realization of WGC. Finally we will present the results of our analysis in Section 6.
@@ II.1 Topological Photon Sphere
To study the photon sphere, we use the topological method instead of the traditional method, a different method that, due to the general characteristics of its mathematical structure 21.1 ; 21.2 , has been extended not only to the study of light rings but also to thermodynamics21.3 ; 21.4 ; 21.5 ; 21.6 ; 21.7 ; 21.8 ; 21.9 ; 21.10 .
In order to use the concept of mapping and topological charge, we need a vector field. If $\varphi_{r}$ and $\varphi_{\theta}$ are components of this general vector field, that is:18
$\varphi=(\varphi_{r},\varphi_{\theta}),$ (1)
also here, we can rewrite the vector as $\varphi=||\varphi||e^{i\theta}$, where $||\varphi||=\sqrt{\varphi_{j}\varphi_{j}}$, or $\varphi=\varphi_{r}+i\varphi_{\theta}$. Based on this, the normalized vector is defined as,
$n^{j}=\frac{\varphi_{j}}{||\varphi||},$ (2)
where $j=1,2$ and $(\varphi_{1}=\varphi_{r})$ , $(\varphi_{2}=\varphi_{\theta})$. In the next stage, with respect to the necessity of spherical symmetry as a prerequisite for studying this method, and given the most general form of the metric in 4-dimensional form we have:
@@
$\begin{split}\mathbb{H}(r,\theta)=\sqrt{\frac{-g_{tt}}{g_{\phi\phi}}}=\frac{1}{\sin\theta}\bigg{(}\frac{f(r)}{h(r)}\bigg{)}^{1/2}.\end{split}$ (4)
With respect to this potential, now the vector field components become as follows:
$\begin{split}&\varphi_{r}=\frac{\partial_{r}\mathbb{H}}{\sqrt{g_{rr}}}=\sqrt{g(r)}\partial_{r}\mathbb{H},\\ &\varphi_{\theta}=\frac{\partial_{\theta}\mathbb{H}}{\sqrt{g_{\theta\theta}}}=\frac{\partial_{\theta}\mathbb{H}}{\sqrt{h(r)}}.\end{split}$ (5)
Now, using the above relations and considering the definition of topological charges, for each photon sphere we can assign a charge. To avoid repetition, a detailed description of how to calculate it is available in references 1 ; 18 ; 19 ; 20 ; 21 .
@@ II.2 Time-like Circular Orbits (TCOs)
Although the geometry and dynamics of spacetime within the event horizon seem beyond our reach, beyond the event horizon, our field equations retain their logical applicability. This allows us to precisely identify and classify geometric-kinematic boundaries. Accordingly, we can consider the following classifications:
1. Light-like Geodesics : Massless particles and photons trajectories.
2. Timelike Geodesics: Massive objects trajectories .
Both types of theses geodesics, depending on the characteristics of the gravitational potential, may either lack observable turning points or possess a turning point, forming loops. These loops can be stable or unstable. In the case of null geodesics, these loops are referred to as photon spheres, while in timelike geodesics, they are known as Timelike Circular Orbits (TCOs).
To better understand the behavior of these orbits and avoid mathematical complexities, we make certain assumptions that do not compromise the generality of our analysis. For instance, in this study, we consider a static and axisymmetric spacetime with $\mathbb{Z}_{2}$ symmetry in a 1+3 dimensional framework. Based on the metric equation Eq. (3), we consider the following quantities 22 :
@@
where the energy and angular momentum are denoted by E, L. Now the Lagrangian can be recast as:
$2\mathfrak{L}=-\frac{\mathbb{A}}{\mathbb{B}}=\varrho,$ (8)
where $\varrho=-1,0$ for time-like, light-like geodesics, respectively. With the above Lagrangian, the effective potential can be rewritten as follows:
@@
For an observer at infinity, the angular velocity in terms of metric parameters will be as follows::
@@
we also consider the $\beta$ quantity in the following form:
@@ III Non-Commutative black hole in 4D Einstein-Gauss-Bonnet theory
The NC parameter is $\Xi$, the total mass of the source is m, $\alpha$ is a coupling constant with dimension of $(length)^{2}$ which is positive in heterotic string theory and the lower incomplete Gamma function is represented with $\gamma$, which can be written in the following form in terms of $\Gamma$ (the upper incomplete Gamma function) :
@@
As previously stated, the smearing of the mass distribution in Non-Commutative Black Holes (NCBH),
@@ III.1 Topological Photon Sphere
$\begin{split}\mathbb{H}=\frac{\sqrt{4+\frac{\left(1-\sqrt{h}\right)r^{2}}{\alpha}}}{2\sin\!\left(\theta\right)r}.\end{split}$ (16)
@@
$\begin{split}\varphi_{r}=-\frac{\left(\varphi_{1}+\varphi_{2}\right)\csc\!\left(\theta\right)}{\pi^{\frac{1}{4}}\sqrt{\frac{16\sqrt{\pi}\,\mathrm{erf}\left(\frac{\sqrt{\frac{r^{2}}{\Xi}}}{2}\right)\alpha m+\sqrt{\pi}\,r^{3}-16\sqrt{\frac{r^{2}}{\Xi}}\,{\mathrm{e}}^{-\frac{r^{2}}{4\Xi}}\alpha m}{r^{3}}}\,r^{3}\Xi}.\end{split}$ (17)
$\begin{split}\varphi_{\theta}=-\frac{\sqrt{4+\frac{\left(1-\sqrt{h}\right)r^{2}}{\alpha}}\,\cos\!\left(\theta\right)}{2\sin\!\left(\theta\right)^{2}r^{2}}.\end{split}$ (18)
@@
Figure 2: Fig (2a): The photon sphere location at $(r,\theta)=(2.736931688316,1.57)$ with respect to $(\Xi=0.001,\alpha=0.39,m=1)$ in the $(r-\theta)$ plane of the normal vector field $n$ , (2b): the topological potential H(r) for NCEGB BH
In the Fig. (2) we can see that the appearance photon sphere has a Total Topological Charge (TTC) of -1 and, This, as can be seen in the H diagram, was equivalent to an unstable maximum, Fig. (2). Consequently, this case presents a black hole containing an unstable photon sphere.
@@
Figure 3: Fig (3a): The photon spheres locations at $(r,\theta)=(2.285600868252,1.57)$ and $(r,\theta)=(2.305792105113,1.57)$ with respect to $(\Xi=0.2488,\alpha=0.39,m=1)$ in the $(r-\theta)$ plane of the normal vector field $n$ , (3b): the topological potential H(r) for NCEGB BH
In the horizonless region, with $\Xi$ = 0.2488, we can see the appearance of two photon spheres with topological charges of -1 and 1, Fig. (3), which indicating a spacetime with TTC=0. This situation is corresponded to the existence of the minimum and maximum in the H diagram, as we can see in the Fig. (3). All of the above statements mean that in this case we will be dealing with a naked singularity.
@@
Table 1: *Unauthorized region: The region with negative or imaginary roots of $\varphi$.
$R_{PLPS}$: the minimum or maximum possible radius for the appearance of an unstable photon sphere.
@@
Table 2: *Unauthorized region: The region with negative or imaginary roots of $\varphi$.
@@ III.2 TCOs
$\mathbb{A}=E^{2}r^{2}+\frac{L^{2}\left(r^{2}\sqrt{\frac{16\sqrt{\pi}\,\mathrm{erf}\left(\frac{\sqrt{\frac{r^{2}}{\Xi}}}{2}\right)\alpha m+\sqrt{\pi}\,r^{3}-16{\mathrm{e}}^{-\frac{r^{2}}{4\Xi}}\sqrt{\frac{r^{2}}{\Xi}}\,\alpha m}{r^{3}}}-\left(r^{2}+4\alpha\right)\pi^{\frac{1}{4}}\right)}{4\pi^{\frac{1}{4}}\alpha}.$ (19)
$\mathbb{B}=-\frac{r^{2}\left(r^{2}\sqrt{\frac{16\sqrt{\pi}\,\mathrm{erf}\left(\frac{\sqrt{\frac{r^{2}}{\Xi}}}{2}\right)\alpha m+\sqrt{\pi}\,r^{3}-16{\mathrm{e}}^{-\frac{r^{2}}{4\Xi}}\sqrt{\frac{r^{2}}{\Xi}}\,\alpha m}{r^{3}}}-\left(r^{2}+4\alpha\right)\pi^{\frac{1}{4}}\right)}{4\pi^{\frac{1}{4}}\alpha}.$ (20)
@@
Figure 4: With respect to $\Xi=0.001,\alpha=0.39,m=1$, Fig (4a): the $\beta$ diagram in the black hole form and (4b): The space classification and MSCO location in the black hole mode
@@
Figure 5: With respect to $\Xi=0.2488,\alpha=0.39,m=1$, Fig (5a): the $\beta$ diagram for structure in naked singularity form , (5b): Enlarging on the forbidden area in the diagram $\beta$ and (5c): The space classification and MSCO location in the naked singularity mode